# The Dirichlet heat trace for domains with curved corners

### v1

@@ 3.4.1. The little trace and the convolution
$\int_{0}^{t}\int_{0}^{\infty}\int_{0}^{\alpha_{0}}(t-s)^{-1}a(\frac{r}{\sqrt{t-s}},\theta,\frac{r^{\prime}}{\sqrt{t-s}},\theta^{\prime})s^{-3/2}b(\frac{r^{\prime}}{\sqrt{s}},\theta^{\prime},\frac{r^{\prime\prime}}{\sqrt{s}},\theta^{\prime\prime})\,r^{\prime}\,d\theta^{\prime}\,dr^{\prime}\,ds.$
@@
(47) $H_{\text{sub}}\cong-T^{-1}\int_{0}^{1}\int_{0}^{\infty}\int_{0}^{\alpha_{0}}(1-\sigma)^{-1}\sigma^{-3/2}\\ \times a\left(\frac{R}{\sqrt{1-\sigma}},\theta,\frac{R^{\prime}}{\sqrt{1-\sigma}},\theta^{\prime}\right)b\left(\frac{R^{\prime}}{\sqrt{\sigma}},\theta^{\prime\prime}\frac{R^{\prime\prime}}{\sqrt{\sigma}},\theta^{\prime\prime}\right)\,R^{\prime}\,d\theta^{\prime}\,dR^{\prime}\,d\sigma.$
@@
(48) $(\textrm{tr }H_{\text{sub}})(R,\theta)=-T^{-1}\int_{0}^{1}\int_{0}^{\infty}\int_{0}^{\alpha_{0}}(1-\sigma)^{-1}\sigma^{-1/2}\\ \times a\left(\frac{R}{\sqrt{1-\sigma}},\theta,\frac{\sqrt{\sigma}\tilde{R}^{\prime}}{\sqrt{1-\sigma}},\theta^{\prime}\right)b\left(\tilde{R}^{\prime},\theta^{\prime},\frac{R}{\sqrt{\sigma}},\theta\right)\,\tilde{R}^{\prime}\,d\theta^{\prime}\,d\tilde{R}^{\prime}\,d\sigma.$
To find the contribution to the heat trace, we first restrict this kernel to the diagonal by setting $(R^{\prime},\theta^{\prime})=(R,\theta)$, then we integrate with respect to $r\,dr\,d\theta=T^{2}R\,dR\,d\theta$. By Melrose’s pushforward theorem (see [MelroseAPS]; see also [GrieserBcalc, Thm 3.10] for a precise statement), the trace contribution is given by a renormalized integral. Writing down this renormalized integral yields (4).
@@ 4. The case $\alpha=\pi/2$
Our goal here is to compute through the $t^{1/2}$ term in the short-time expansion of the heat trace of the Laplacian on a domain $\Omega$ which is a half-disk of radius 1, with Dirichlet boundary conditions. This heat trace is denoted
@@
The heat trace $\mathrm{Tr}\ e^{-t\Delta_{D}}$ is well understood and has a full polyhomogeneous asymptotic expansion in half-integer powers as $t\to 0$. The beginning of this expansion is, from [EGS_2017],
@@
For $H(t)$, we use the asymptotics of Bessel function zeroes due to McMahon [NIST:DLMF, (10.21.19)]:
@@
We could reverse engineer $c$ from the known expansion from [NRS] but there is no need.
We are now in a position to state what the contribution of each corner of the half-disk to the $t^{1/2}$ term is. The $t^{1/2}$ term above is the sum of a curvature integral over the boundary and a pair of identical corner contributions. From [EGS_2017] we see that the $\frac{\sqrt{\pi}}{256}$ is precisely the curvature integral. Dividing the remainder in half shows
@@
It remains only to prove the Lemmas.
@@ 4.1. Proofs of the Lemmas
We know from [NRS] that $\mathrm{Tr}\ e^{-t\Delta_{\Omega}}$ has a full polyhomogeneous asymptotic expansion, and therefore by (51), so does $H(t)$. By Lemma 3, $\tilde{H}(t)$ has at least a polyhomogeneous expansion up to $O(t)$, and therefore so does $\hat{H}(t)$. The upshot of this is that one may define a zeta function
@@
where $\zeta(s,a)$ is the Hurwitz zeta function. The function $\zeta(s,a)$ has a single, simple pole at $s=1$, with residue 1 [NIST:DLMF, 25.11(i)]. So $\hat{\zeta}(s)$ has a single, simple pole at $s=1/2$, and its residue there is $\frac{1}{2\pi}$. Thus the function

### v2

@@ 3.4.1. The little trace and the convolution
$\int_{0}^{t}\int_{0}^{\infty}\int_{0}^{\alpha}(t-s)^{-1}a(\frac{r}{\sqrt{t-s}},\theta,\frac{r^{\prime}}{\sqrt{t-s}},\theta^{\prime})s^{-3/2}b(\frac{r^{\prime}}{\sqrt{s}},\theta^{\prime},\frac{r^{\prime\prime}}{\sqrt{s}},\theta^{\prime\prime})\,r^{\prime}\,d\theta^{\prime}\,dr^{\prime}\,ds.$
@@
(59) $H_{\text{sub}}\cong-T^{-1}\int_{0}^{1}\int_{0}^{\infty}\int_{0}^{\alpha}(1-\sigma)^{-1}\sigma^{-3/2}\\ \times a\left(\frac{R}{\sqrt{1-\sigma}},\theta,\frac{R^{\prime}}{\sqrt{1-\sigma}},\theta^{\prime}\right)b\left(\frac{R^{\prime}}{\sqrt{\sigma}},\theta^{\prime},\frac{R^{\prime\prime}}{\sqrt{\sigma}},\theta^{\prime\prime}\right)\,R^{\prime}\,d\theta^{\prime}\,dR^{\prime}\,d\sigma.$
@@
(60) $(\textrm{tr }H_{\text{sub}})(R,\theta)=-T^{-1}\int_{0}^{1}\int_{0}^{\infty}\int_{0}^{\alpha}(1-\sigma)^{-1}\sigma^{-1/2}\\ \times a\left(\frac{R}{\sqrt{1-\sigma}},\theta,\frac{\sqrt{\sigma}\tilde{R}^{\prime}}{\sqrt{1-\sigma}},\theta^{\prime}\right)b\left(\tilde{R}^{\prime},\theta^{\prime},\frac{R}{\sqrt{\sigma}},\theta\right)\,\tilde{R}^{\prime}\,d\theta^{\prime}\,d\tilde{R}^{\prime}\,d\sigma.$
To find the contribution to the heat trace, we first restrict this kernel to the diagonal by setting $(R^{\prime},\theta^{\prime})=(R,\theta)$, then we integrate with respect to $r\,dr\,d\theta=T^{2}R\,dR\,d\theta$. By Melrose’s pushforward theorem (see [9]; see also [5, Theorem 3.10] for a precise statement), the trace contribution is given by a renormalized integral. Writing down this renormalized integral yields (4).
The last thing to do is to address the dependence of this expression on $\kappa_{\pm}$. By (58), the kernel $b$ does appear to depend on $\theta_{0}$ through factors such as $\cos(\theta+\theta_{0})$. However, integrating in the angular variables removes this dependence. Let $I(\theta_{0})$ denote the right-hand side of (4). From (58), the $\theta_{0}$ dependence enters only through shifted angles of the form $\theta+\theta_{0}$ and $\phi+\theta_{0}$, while the remaining angular dependence, coming from the $D^{2}$ term, occurs through differences such as $\theta-\phi$. Differentiating under the integral sign therefore yields
$\frac{d}{d\theta_{0}}I(\theta_{0})=\iint_{(0,\alpha)\times(0,\alpha)}(\partial_{\theta}+\partial_{\phi})(\cdots)\,d\theta\,d\phi,$
where $(\cdots)$ denotes the $\theta,\phi$ dependent integrand in (4). Integrating by parts in $\theta$ and $\phi$ reduces this to boundary terms at $\theta,\phi\in\{0,\alpha\}$, and these vanish for Dirichlet boundary conditions since the sector kernel $a$ in (43) contains the factors $\varphi_{j}(\theta)$ and $\varphi_{j}(\phi)$ with $\varphi_{j}(0)=\varphi_{j}(\alpha)=0$. Hence $\frac{d}{d\theta_{0}}I(\theta_{0})=0$, so $c_{1/2}(\alpha)=I(\theta_{0})$ is independent of $\theta_{0}$. This justifies writing $c_{1/2}(\alpha)$, rather than $c_{1/2}(\alpha,\kappa_{-},\kappa_{+})$, on the left-hand side of (4).
@@ 4. The case $\alpha=\pi/2$
4.1. The Dirichlet case
Our goal here is to compute through the $t^{1/2}$ term in the short-time expansion of the heat trace of the Laplacian on a domain $\Omega$ which is a half-disk of radius 1, with Dirichlet boundary conditions. This provides a concrete test case in which the corner term can be read off from an explicit spectrum and compared directly with (1) and (3). This heat trace is denoted
@@ 4.1. The Dirichlet case
The heat trace $\mathrm{Tr}\ e^{-t\Delta_{D}}$ is well understood and has a full polyhomogeneous asymptotic expansion in half-integer powers as $t\to 0$. The beginning of this expansion is, from [4],
@@
For $H(t)$, we use the asymptotics of Bessel function zeroes due to McMahon [3, (10.21.19)]:
@@
We could reverse engineer $c$ from the known expansion from [10] but there is no need.
We are now in a position to state what the contribution of each corner of the half-disk to the $t^{1/2}$ term is. The $t^{1/2}$ term above is the sum of a curvature integral over the boundary and a pair of identical corner contributions. From [4] we see that the $\frac{\sqrt{\pi}}{256}$ is precisely the curvature integral. Dividing the remainder in half shows
@@
Combined with (3), and the fact that $r_{0}(\frac{\pi}{2},1,0)=\frac{1}{2}$, this identifies $c_{1/2}(\pi/2)$ and confirms that right-angled curved corners contribute nontrivially at order $t^{1/2}$. It remains only to prove the Lemmas.
@@ 4.2. Proofs of the Lemmas
We know from [10] that $\mathrm{Tr}\ e^{-t\Delta_{\Omega}}$ has a full polyhomogeneous asymptotic expansion, and therefore by (63), so does $H(t)$. By Lemma 3, $\tilde{H}(t)$ has at least a polyhomogeneous expansion up to $O(t)$, and therefore so does $\hat{H}(t)$. The upshot of this is that one may define a zeta function
@@
where $\zeta(s,a)$ is the Hurwitz zeta function. The function $\zeta(s,a)$ has a single, simple pole at $s=1$, with residue 1 [3, 25.11(i)]. So $\hat{\zeta}(s)$ has a single, simple pole at $s=1/2$, and its residue there is $\frac{1}{2\pi}$. Thus the function
@@
4.3. A Neumann right-angle corner contribution
This section computes the order-$t^{1/2}$ contribution coming from the two right-angle corners of the unit half-disk with Neumann boundary conditions. The computation uses the explicit Neumann spectrum on the disk and half-disk, together with the smooth-boundary coefficient for Neumann boundary conditions. The final outcome is an explicit value for the corner remainder in this model geometry, which we record as $\mathcal{C}^{N}_{1/2}\bigl(\frac{\pi}{2},1,0\bigr)$.
Let $\Omega\subset\mathbb{R}^{2}$ be the unit half-disk
$\Omega=\{(r,\theta):0<r<1,\ 0<\theta<\pi\},$
endowed with Neumann boundary conditions along both the diameter and the semicircle. Let $D$ denote the unit disk, also with Neumann boundary conditions, and write $\Delta^{N}_{\Omega}$ and $\Delta^{N}_{D}$ for the corresponding Laplacians. For the unit disk with Neumann boundary conditions, the smooth-boundary heat trace expansion gives (see Proposition 8 of [4])
(82) $\mathrm{Tr}\ e^{-t\Delta^{N}_{D}}=\frac{1}{4}t^{-1}+\frac{\sqrt{\pi}}{4}t^{-1/2}+\frac{1}{6}+\frac{5\sqrt{\pi}}{128}t^{1/2}+O(t).$
Separation of variables shows that the Neumann spectrum on the disk is
$\mathrm{spec}(\Delta_{D}^{N})=\{0\}\ \cup\ \{(j^{\prime}_{m,k})^{2}:\ m\in\mathbb{N}_{0},\ k\in\mathbb{N}\},$
where $j^{\prime}_{m,k}$ is the $k$th positive zero of $J_{m}^{\prime}$, with multiplicity $1$ for $m=0$ and multiplicity $2$ for $m\geq 1$ (the $\cos(m\theta)$ and $\sin(m\theta)$ modes). On the half-disk, the Neumann condition on $\theta=0,\pi$ forces $\partial_{\theta}u=0$, hence only the $\cos(m\theta)$ modes occur, so each $(j^{\prime}_{m,k})^{2}$ has multiplicity $1$. Thus
(83) $\mathrm{Tr}\ e^{-t\Delta^{N}_{\Omega}}=\frac{1}{2}\,\mathrm{Tr}\ e^{-t\Delta^{N}_{D}}+\frac{1}{2}\,H^{N}(t),$
where the correction term is precisely the $m=0$ contribution
$H^{N}(t):=\sum_{\lambda\in\{0\}\cup\{(j^{\prime}_{0,k})^{2}\}_{k\geq 1}}e^{-t\lambda}=1+\sum_{k=1}^{\infty}e^{-t(j^{\prime}_{0,k})^{2}}.$
Using $J_{0}^{\prime}(x)=-J_{1}(x)$, this may be rewritten as
(84) $H^{N}(t)=1+\sum_{k=1}^{\infty}e^{-tj_{1,k}^{2}},$
where $\{j_{1,k}\}_{k\geq 1}$ are the positive zeros of $J_{1}$.
McMahon’s asymptotics for Bessel zeros (cf. [3, (10.21.19)]) give
(85) $j_{1,k}=\Bigl(k+\frac{1}{4}\Bigr)\pi-\frac{3}{8\bigl(k+\frac{1}{4}\bigr)\pi}+O(k^{-3}),\qquad k\to\infty,$
hence
(86) $j_{1,k}^{2}=\Bigl(k+\frac{1}{4}\Bigr)^{2}\pi^{2}-\frac{3}{4}+O(k^{-2}).$
Define the model series
(87) $\widetilde{H}^{N}(t):=1+\sum_{k=1}^{\infty}\exp\Bigl[-t\Bigl(\Bigl(k+\frac{1}{4}\Bigr)^{2}\pi^{2}-\frac{3}{4}\Bigr)\Bigr].$
Lemma 5.
As $t\to 0$ one has $H^{N}(t)-\widetilde{H}^{N}(t)=O(t)$.
Proof.
This is identical to the proof of Lemma 3. ∎
Lemma 6.
As $t\to 0$,
(88) $\widetilde{H}^{N}(t)=\frac{1}{2\sqrt{\pi}}\,t^{-1/2}+c_{N}+\frac{3}{8\sqrt{\pi}}\,t^{1/2}+O(t),$
for some constant $c_{N}\in\mathbb{R}$.
Proof.
Write $\widetilde{H}^{N}(t)=e^{3t/4}\,\widehat{H}^{N}(t)$ with
$\widehat{H}^{N}(t):=e^{-3t/4}\widetilde{H}^{N}(t)=e^{-3t/4}+\sum_{k=1}^{\infty}\exp\Bigl[-t\Bigl(k+\frac{1}{4}\Bigr)^{2}\pi^{2}\Bigr].$
As in the proof of Lemma 4, the existence of a polyhomogeneous expansion for $\widehat{H}^{N}(t)$ through $O(t)$ allows one to define the Mellin transform
$\widehat{\zeta}_{N}(s)=\frac{1}{\Gamma(s)}\int_{0}^{\infty}\widehat{H}^{N}(t)\,t^{s-1}\,dt=\sum_{k=0}^{\infty}\bigl(\pi(k+\tfrac{5}{4})\bigr)^{-2s}=\pi^{-2s}\zeta(2s,\tfrac{5}{4}),$
initially for $\Re(s)>>0$ and then by meromorphic continuation.
Since $\zeta(2s,\tfrac{5}{4})$ has a simple pole at $2s=1$ with residue $1$, the function $\widehat{\zeta}_{N}$ has a simple pole at $s=\tfrac{1}{2}$ with residue $1/(2\pi)$. It follows that $\int_{0}^{\infty}\widehat{H}^{N}(t)t^{s-1}\,dt=\Gamma(s)\widehat{\zeta}_{N}(s)$ has residue $\Gamma(1/2)/(2\pi)=1/(2\sqrt{\pi})$ at $s=1/2$, and hence
$\widehat{H}^{N}(t)=\frac{1}{2\sqrt{\pi}}\,t^{-1/2}+c_{N}+O(t).$
Multiplying by $e^{3t/4}=1+\frac{3}{4}t+O(t^{2})$ yields (88). ∎
Combining Lemmas 5 and 6 shows that $H^{N}(t)$ itself satisfies (88) (with a possibly different constant term), up to $O(t)$. Substituting (82) and (88) into (83) gives the half-disk expansion
(89) $\mathrm{Tr}\ e^{-t\Delta^{N}_{\Omega}}=\frac{1}{8}\,t^{-1}+\Bigl(\frac{\sqrt{\pi}}{8}+\frac{1}{4\sqrt{\pi}}\Bigr)t^{-1/2}+\text{\emph{(const)}}+\Bigl(\frac{5\sqrt{\pi}}{256}+\frac{3}{16\sqrt{\pi}}\Bigr)t^{1/2}+O(t).$
The coefficient of $t^{1/2}$ in (89) may be split into a smooth-boundary term and a corner remainder by subtracting the smooth-boundary contribution for Neumann boundary conditions. For the unit disk, $\int_{\partial D}\kappa^{2}\,ds=2\pi$, and the $t^{1/2}$ coefficient $\frac{5\sqrt{\pi}}{128}$ in (82) corresponds to the smooth-boundary constant $\frac{5}{256\sqrt{\pi}}$. For the half-disk, $\partial\Omega$ consists of a semicircle of curvature $1$ and length $\pi$ and a diameter of curvature 0, so
$\frac{5}{256\sqrt{\pi}}\int_{\partial\Omega}\kappa^{2}\,ds=\frac{5}{256\sqrt{\pi}}\cdot\pi=\frac{5\sqrt{\pi}}{256}.$
Subtracting this from the $t^{1/2}$ coefficient in (89) leaves a remainder $\frac{3}{16\sqrt{\pi}}$. The two endpoints of the semicircle are congruent corners, so they contribute equally. This motivates the definition
(90) $\mathcal{C}^{N}_{1/2}\Bigl(\frac{\pi}{2},1,0\Bigr):=\frac{1}{2}\cdot\frac{3}{16\sqrt{\pi}}=\frac{3}{32\sqrt{\pi}}.$
Here $(\kappa_{+},\kappa_{-})=(1,0)$ records that one incident boundary arc has curvature $1$ (the semicircle) and the other has curvature 0 (the diameter).
Remark.
Equation (90) records the corner value $\mathcal{C}^{N}_{1/2}\bigl(\frac{\pi}{2},1,0\bigr)=\frac{3}{32\sqrt{\pi}}$. If a Neumann factorization of the form
$\mathcal{C}^{N}_{1/2}(\alpha,\kappa_{+},\kappa_{-})=c^{N}_{1/2}(\alpha)\,r_{0}(\alpha,\kappa_{+},\kappa_{-})$
holds for $\alpha=\pi/2$, then since $r_{0}(\pi/2,1,0)=\frac{1}{2}$ by (10), one obtains
$c^{N}_{1/2}\Bigl(\frac{\pi}{2}\Bigr)=\frac{\mathcal{C}^{N}_{1/2}\bigl(\frac{\pi}{2},1,0\bigr)}{r_{0}(\pi/2,1,0)}=\frac{3}{16\sqrt{\pi}}.$